# SPARKE: Scalable Prompt-Aware Diversity Guidance in Diffusion Models via RKE Score

### v1

@@ Appendix C Implementation Details and Hyperparameters
In the kernel-based guidance experiments of SPARKE and the baselines with kernel entropy diversity scores, we considered a Gaussian kernel, which consistently led to higher output scores in comparison to the other standard cosine similarity kernel (see Section D for ablation studies). We used the same Gaussian kernel bandwidth $\sigma$ in the RKE and Vendi experiments, and the bandwidth parameter choice matches the selected value in [15, 13]. The numerical experiments were conducted on 4$\times$×NVIDIA GeForce RTX 4090 GPUs, each of which has 22.5 GB of memory.
@@ C.1 Experimental Configuration for Table 2
We evaluated the methods listed in Table 2 in the following setting. We used Stable Diffusion 2.1 with a resolution of 1024$\times$1024, a fixed classifier-free guidance scale of $W_{\text{CFG}}=7.5$, and 50 inference steps using the DPM solver. We used the first 10,000 prompts of the MS-COCO 2014 validation set and fixed the generation seed to be able to compare the effect of the methods. For the methods, we used the following configuration to generate the results reported in Table 2.
The hyperparameter tuning was performed by performing cross-validation on the in-batch similarity score, selecting the hyperparameter values that optimized this alignment-based metric. Note that the in-batch similarity score accounts for both text-image consistency and inter-sample diversity as discussed in [49]. Due to GPU memory requirements (as mentioned in Section B.2 in [49]), we were unable to evaluate the Particle Guide baseline [49] on SD v2.1. Following the provided implementation of this baseline for SD v1.5, we conducted this baseline’s experiments only on SD v1.5.
@@
CADS. Following the discussion in Table 13 of [14], we set the threshold parameters as $\tau_{1}=0.6$ and $\tau_{2}=0.9$, $\psi=1$, and used a noise scale of 0.25.
c-VSG. We note that the reference [10] considered GeoDE [62] and DollarStreet [70] datasets, in which multiple samples exist per input prompt. On the other hand, in our experiments, we considered the standard MSCOCO prompt set where for each prompt corresponds we access a single image, making the contextualized Vendi guidance baseline in [10] not directly applicable. Therefore, we simulated the non-contextualized version of VSG. For selecting the Vendui score guidance scale, we performed validation over the set $\{0,0.04,0.05,0.06,0.07\}$, following the procedure in [10]. A guidance frequency of 5 was used, consistent with the original implementation. To maintain stable gradient computation for the Vendi score, we implemented a sliding window of 150 most recently generated samples, as gradient calculations became numerically unstable for some steps beyond this threshold.
latent RKE Guidance. We used a Gaussian kernel with bandwidth $\sigma_{img}=0.8$ and used $\eta=0.03$ as the weight of RKE guidance. To balance the effects of the diversity guidance in sample generation, the RKE guidance update was applied every 10 reverse-diffusion steps in the diffusion process, which is similar to the implementation of Vendi score guidance in [10]. Unlike VSG, which requires a sample window limit, RKE guidance operates without window size constraints, leveraging the complete history of generated samples for gradient computation.
@@ C.2 Experiment Settings in the results of Table 3 and Figure 3
We conducted additional experiments using Stable Diffusion v1.5 with a resolution of 512$\times$512. We created a prompt set by performing K-Means clustering on the MSCOCO 2014 validation prompts. Specifically, we clustered the MS-COCO prompt dataset via the spectral clustering in the CLIP embedding space, into 40 groups and randomly drawn 50 prompts from each cluster, resulting in a total of 2,000 prompts. To evaluate the performance of our method with the baselines with different seeds, we generated five samples per prompt using seeds 0, 1, 2, 3, and 4, yielding a total of 10,000 images for each method. The complete list of prompts is included in the supplementary materials.
@@
Particle Guide. We applied the original implementation provided by the repository of [49]. Since the method operates on repeated generations of the same prompt, we generated five samples per prompt using seeds 0 through 4. The coefficient parameter was set to 30, following the original implementation.
@@ Appendix D Additional Numerical Results
Effect of Classifier-Free Guidance scale on SPARKE guidance. We study the Classifier-free guidance’s impact on SPARKE for the Stable Diffusion 1.5 model, demonstrating how our method alleviates the known quality-diversity trade-off [46]. We changed the CFG scale from 2 to 8 while using the same setting of SD 1.5 as mentioned in Section C.2. We evaluated SPARKE’s effect on the quality-diversity trade-off using Precision/Recall metrics in Figure 14 and Density/Coverage metrics in Figure 13. Our results show that SPARKE guidance maintains diversity in a high CFG scale while preserving acceptable generation quality. Additionally, the Vendi score and Kernel Distance (KD) evaluations in Figure 12 show the same trend. Qualitative comparisons of generated samples at CFG scales $w=4,6,8$ are presented in Figure 15, demonstrating that SPARKE produces diverse outputs without compromising quality.
@@
Additional quantitative comparison with the baseline diversity-guided diffusion-based sample generations. Similar to Table 2, we compare different guidance methods on Stable Diffusion 1.5 [53] in Table 3.
@@
SD v1.5 [53] 30.15 1.045 73.86 25.41 350.28 81.25
c-VSG [10] 27.80 1.078 74.92 26.78 357.39 79.71
CADS [14] 28.94 1.049 75.44 27.60 360.99 78.23
Particle Guide [49] 29.50 1.056 72.58 27.24 345.50 79.50
@@
Latent space vs. ambient space. We compared latent-space guidance (as in SPARKE) with the CLIP-embedded ambient-space guidance in c-VSG [10] using three categories from the GeoDE dataset [62] and modified it with GPT-4o [63]. We use a similar prompt template as mentioned in Table 7 in [10] to add more details to the prompt. We provide additional samples in Figure 16. We use the following template:
@@
Additional qualitative comparisons of vanilla, baseline diversity guided, and SPARKE prompt-aware diversity guided diffusion models. We compared the SPARKE prompt-aware diversity guidance method with state-of-the-art conditional latent diffusion models, including Stable Diffusion 2.1, Stable Diffusion XL [5] and PixArt-$\Sigma$ [71].
Additional qualitative comparisons between prompt-aware (conditional RKE score) and prompt-unaware (RKE score) diversity guidance methods are presented on SD 2.1 (Figure 17). Building on these observations, our results demonstrate that SPARKE improves prompt-aware diversity more effectively than prompt-unaware RKE guidance, as quantified by In-batch Similarity [49].
@@
Pose to image generation. We further extended our evaluation to pose-to-image conditional diffusion models, which operate under different conditions compared to text-to-image models. In Figure 20, SPARKE (conditional RKE guidance) was compared with unconditional RKE guidance for pose-to-image generation tasks, implemented with ControlNet and the Stable Diffusion 1.5 backbone. The qualitative results show SPARKE’s efficacy in enhancing the diversity of generated images when extended to different conditions.
Figure 20: Comparison of ControlNet Model vs. ControlNet guided by SPARKE method.

### v2

@@ Appendix C Implementation Details and Hyperparameters
[…]
@@ C.1 Experimental Configuration for Table 1
We evaluated the methods listed in Table 1 in the following setting. We used Stable Diffusion 2.1 with a resolution of $\times1024$, a fixed classifier-free guidance scale of $W_{\text{CFG}}=7.5$, and 50 inference steps using the DPM solver. We used the first 10,000 prompts of the MS-COCO 2014 validation set and fixed the generation seed to be able to compare the effect of the methods. For the methods, we used the following configuration to generate the results reported in Table 1.
The hyperparameter tuning was performed by performing cross-validation on the in-batch similarity score, selecting the hyperparameter values that optimized this alignment-based metric. Note that the in-batch similarity score accounts for both text-image consistency and inter-sample diversity as discussed in corso2024particleguidance . Due to GPU memory requirements (as mentioned in Section B.2 in corso2024particleguidance ), we were unable to evaluate the Particle Guide baseline corso2024particleguidance on SD v2.1. Following the provided implementation of this baseline for SD v1.5, we conducted this baseline’s experiments only on SD v1.5.
@@
CADS. Following the discussion in Table 13 of sadat2024cads , we set the threshold parameters as $\tau_{1}=0.6$ and $\tau_{2}=0.9$, $\psi=1$, and used a noise scale of 0.25.
c-VSG. We note that the reference askari2024improving considered GeoDE ramaswamy2022geode and DollarStreet dollar_street datasets, in which multiple samples exist per input prompt. On the other hand, in our experiments, we considered the standard MSCOCO prompt set where for each prompt corresponds we access a single image, making the contextualized Vendi guidance baseline in askari2024improving not directly applicable. Therefore, we simulated the non-contextualized version of VSG. For selecting the Vendui score guidance scale, we performed validation over the set $\{0,0.04,0.05,0.06,0.07\}$, following the procedure in askari2024improving . A guidance frequency of 5 was used, consistent with the original implementation. To maintain stable gradient computation for the Vendi score, we implemented a sliding window of 150 most recently generated samples, as gradient calculations became numerically unstable for some steps beyond this threshold.
latent RKE Guidance. We used a Gaussian kernel with bandwidth $\sigma_{img}=0.8$ and used $\eta=0.03$ as the weight of RKE guidance. To balance the effects of the diversity guidance in sample generation, the RKE guidance update was applied every 10 reverse-diffusion steps in the diffusion process, which is similar to the implementation of Vendi score guidance in askari2024improving . Unlike VSG, which requires a sample window limit, RKE guidance operates without window size constraints, leveraging the complete history of generated samples for gradient computation.
@@ C.2 Experiment Settings in the results of Table 4 and Figure 4
We conducted additional experiments using Stable Diffusion v1.5 with a resolution of $\times512$. We created a prompt set by performing K-Means clustering on the MSCOCO 2014 validation prompts. Specifically, we clustered the MS-COCO prompt dataset via the spectral clustering in the CLIP embedding space, into 40 groups and randomly drawn 50 prompts from each cluster, resulting in a total of 2,000 prompts. To evaluate the performance of our method with the baselines with different seeds, we generated five samples per prompt using seeds 0, 1, 2, 3, and 4, yielding a total of 10,000 images for each method. The complete list of prompts is included in the supplementary materials.
@@
Particle Guide. We applied the original implementation provided by the repository of corso2024particleguidance . Since the method operates on repeated generations of the same prompt, we generated five samples per prompt using seeds 0 through 4. The coefficient parameter was set to 30, following the original implementation.
@@ Appendix D Additional Numerical Results
Effect of Classifier-Free Guidance scale on SPARKE guidance. We study the Classifier-free guidance’s impact on SPARKE for the Stable Diffusion 1.5 model, demonstrating how our method alleviates the known quality-diversity trade-off ho2022classifier . We changed the CFG scale from 2 to 8 while using the same setting of SD 1.5 as mentioned in Section C.2. We evaluated SPARKE’s effect on the quality-diversity trade-off using Precision/Recall metrics in Figure 14 and Density/Coverage metrics in Figure 13. Our results show that SPARKE guidance maintains diversity in a high CFG scale while preserving acceptable generation quality. Additionally, the Vendi score and Kernel Distance (KD) evaluations in Figure 12 show the same trend. Qualitative comparisons of generated samples at CFG scales $w=4,6,8$ are presented in Figure 15, demonstrating that SPARKE produces diverse outputs without compromising quality.
@@
Additional quantitative comparison with the baseline diversity-guided diffusion-based sample generations. Similar to Table 1, we compare different guidance methods on Stable Diffusion 1.5 Rombach_2022_CVPR in Table 4.
@@
SD v1.5 Rombach_2022_CVPR 30.15 1.045 73.86 25.41 350.28 81.25
c-VSG askari2024improving 27.80 1.078 74.92 26.78 357.39 79.71
CADS sadat2024cads 28.94 1.049 75.44 27.60 360.99 78.23
Particle Guide corso2024particleguidance 29.50 1.056 72.58 27.24 345.50 79.50
@@
Latent space vs. ambient space. We compared latent-space guidance (as in SPARKE) with the CLIP-embedded ambient-space guidance in c-VSG askari2024improving using three categories from the GeoDE dataset ramaswamy2022geode and modified it with GPT-4o openai_gpt-4_2024 . We use a similar prompt template as mentioned in Table 7 in askari2024improving to add more details to the prompt. We provide additional samples in Figure 16. We also provide quantitative results in Table 5 shows that although CLIP-based Vendi scores were slightly higher for CLIP-embedded ambient guidance, which could be due to optimizing Vendi with the CLIP features, the DINOv2-based Vendi scores and in-batch similarity scores indicate higher diversity for the latent Vendi entropy guidance that we proposed for LDMs. Furthermore, latent guidance can be performed with a considerably lower GPU memory (in our implementation $\approx$20GB for the latent case vs. approx\ approx35 GB for the ambient case), resulting in higher computational efficiency.
[…]
(CLIP-embedded) Ambient VSG askari2024improving 5.45 5.34 13.33 0.28
Latent VSG 7.24 4.42 25.22 0.15
We use the following template:
@@
Additional qualitative comparisons of vanilla, baseline diversity guided, and SPARKE prompt-aware diversity guided diffusion models. We compared the SPARKE prompt-aware diversity guidance method with state-of-the-art conditional latent diffusion models, including Stable Diffusion 2.1, Stable Diffusion XL podell2024sdxl and PixArt-$\Sigma$ pixart_sigma .
Additional qualitative comparisons between prompt-aware (conditional RKE score) and prompt-unaware (RKE score) diversity guidance methods are presented on SD 2.1 (Figure 17). Building on these observations, our results demonstrate that SPARKE improves prompt-aware diversity more effectively than prompt-unaware RKE guidance, as quantified by In-batch Similarity corso2024particleguidance .